# Multimodal Datasets with Controllable Mutual Information

### v1

@@ 5 Related Work
The relevance of mutual information to self-supervised learning, particularly contrastive learning, has been studied extensively. For example, the InfoNCE family of contrastive loss functions can be interpreted as bounds on the mutual information between representations [4, 6, 7]. A range of work inspired by the InfoMax principle [36] has argued that the MI between inputs and learned representations is an implicit target for multiview contrastive learning [4, 37, 38, 39]. However, Tschannen et al. [40] find that maximizing the MI alone is not sufficient for learning representations that are useful for downstream tasks, stressing that the relation between estimated MI and representation quality depends strongly on both architecture choice and the form of the mutual information estimator used. More recent work [41, 42, 43, 44] explores this question in further depth. Our ability to generate realistic, complex datasets with known mutual information may enable further progress in determining the role of information maximization in self-supervised learning.
@@
Estimating mutual information from samples is challenging [14], especially in compelling real-world datasets and applications [45, 46, 47]. A rich body of literature covers a variety of mutual information estimation methods, ranging from traditional approaches based on histogram density or $k$-nearest neighbors [16, 17, 19] to neural estimators based on variational approaches [20, 21] or generative modeling [48, 22]. However, benchmarking the efficacy of these estimators on realistic datasets is entirely nontrivial. Most existing approaches are validated on multivariate normal distributions where mutual information is easily controllable, while recent work has explored simple transformations of these distributions to emulate properties of real data [15, 26, 27]. Our work replaces these simple transformations with a flexible bijective mapping learned through flow-based generative modeling [49], enabling construction of highly realistic datasets with analytically tractable mutual information.
@@
Several generative models satisfy the mutual information preservation condition. Discrete-time normalizing flows with coupling layers [32, 33] and Invertible Residual Networks [50] were among the first invertible (bijective) deep generative models. More recently, TarFlow and StarFlow [51, 52] have achieved very strong results on high resolution image generation. The ability of flow-based models to preserve mutual information has been employed in a variety of contexts, including mutual information estimation [22] and developing alternative prescriptions for training flow models [53]. However, employing this property to develop realistic datasets with known mutual information remains a novel contribution of this work.
@@ 6 Conclusion
We present a new framework for generating realistic datasets with many modalities that are designed with known and controllable mutual information. Our dataset generation framework uses interpretable causal models with linear structural equations to construct correlated, normally-distributed latent variables with known mutual information. Blocks of components of these random variables are then fed into invertible (bijective) transformations that map the latent inputs into a realistic feature space while preserving the mutual information content. These realistic and nontrivial datasets enable numerous studies, including benchmarking studies of mutual information estimators. Critically, these datasets will be important for understanding and validating the role of mutual information in various multimodal self-supervised learning strategies, particularly as the number of modalities grows.
@@
To encourage reproducibility, we submit all code as supplementary material. Additionally, we describe the details of our example datasets, including hyperparameters chosen.
Acknowledgments
MP gratefully acknowledges the Center for Computational Astrophysics at the Flatiron Institute for hospitality while a portion of this work was carried out. GM and KC are supported in part by the U.S. Department of Energy (DOE) under Award No. DE-FOA-0002705, KA/OR55/22 (AIHEP).
Appendix A Appendix
A.1 Analytic formulae for covariance matrices and mutual information

### v2

@@ 5 Related Work
The relevance of mutual information (MI) to self-supervised learning (SSL), particularly contrastive learning, has been studied extensively. For example, the InfoNCE family of contrastive loss functions can be interpreted as bounds on the MI between representations [oord2018representation, he2020momentum, caron2020unsupervised]. A range of work inspired by the InfoMax principle [Linsker] has argued that the MI between inputs and learned representations is an implicit target for multiview contrastive learning [oord2018representation, hjelm2019, henaff2020, tsai2021]. However, tschannen2020 find that maximizing the MI alone is not sufficient for learning representations that are useful for downstream tasks, stressing that the relation between estimated MI and representation quality depends strongly on both architecture choice and the form of the MI estimator used. More recent work [tian2020, wang2022, wang2022hypersphere, rodríguezgálvez2023] explores this question in further depth. Our ability to generate realistic, complex datasets with known MI may enable further progress in determining the role of information maximization in self-supervised learning.
@@
Estimating mutual information from samples is challenging [McAllester-2020-FormalLimitations], especially in compelling real-world datasets and applications [holmes:19:estimation, gao:15:strongly-dependent, gao:17:mixture]. A rich body of literature covers a variety of mutual information estimation methods, ranging from traditional approaches based on histogram density or $k$-nearest neighbors [pizer1987adaptive, kozachenko:87:firstnn, kraskov:04:ksg] to neural estimators based on variational approaches [belghazi:18:mine, song:20:understanding] or generative modeling [ao:22:entropy-flows, butakov2024mutual]. However, benchmarking the efficacy of these estimators on realistic datasets is entirely nontrivial. Most existing approaches are validated on multivariate normal distributions where mutual information is easily controllable, while recent work has explored simple transformations of these distributions to emulate properties of real data [czyz2023beyond, czyz2023properties, butakov2023information]. Our work replaces these simple transformations with a flexible bijective mapping learned through flow-based generative modeling [NEURIPS2018_69386f6b], enabling construction of highly realistic datasets with analytically tractable MI.
@@
Several generative models satisfy the mutual information preservation condition. Discrete-time normalizing flows with coupling layers [rezende2015variational, papamakarios2021normalizing] and Invertible Residual Networks [pmlr-v97-behrmann19a] were among the first invertible (bijective) deep generative models. More recently, TarFlow and StarFlow [zhai2024normalizing, gu2025starflow] have achieved very strong results on high resolution image generation. The ability of flow-based models to preserve mutual information has been employed in a variety of contexts, including mutual information estimation [butakov2024mutual] and developing alternative prescriptions for training flow models [ardizzone2021trainingnormalizingflowsinformation]. However, employing this property to develop realistic datasets with known mutual information remains a novel contribution of this work.
@@ 6 Conclusion
We present a new framework for generating realistic datasets with many modalities that are designed with known and controllable mutual information. Our dataset generation framework uses interpretable causal models with linear structural equations to construct correlated, normally-distributed latent variables with known mutual information. Blocks of components of these random variables are then fed into invertible (bijective) transformations that map the latent inputs into a realistic feature space while preserving the mutual information content. These realistic and nontrivial datasets enable numerous studies, including benchmarking studies of mutual information estimators. Critically, these datasets will be important for understanding the role of mutual information in various multimodal self-supervised learning strategies, particularly as the number of modalities grows.
@@
To encourage reproducibility, we submit all code as supplementary material. Additionally, we describe the details of our example datasets, including hyperparameters chosen. Upon publication, we will release the code and sample datasets publicly.
Appendix A Prediction Error for the $\theta$ Regression Task
In Figure 7, we show the distribution of prediction errors $\theta-\hat{\theta}$ for two illustrative points along our MI distribution. The dataset with lower MI (in blue) exhibits a wider spread in errors, indicating decreased regression performance, while the dataset with higher MI (in orange) shows a sharper peak, indicating better regression performance. Moreover, both distributions are centered at 0, indicating that they are not biased in their estimations of $\theta$.
Figure 7: An example distribution of prediction error $\theta-\hat{\theta}$ for two representative MI values.
Appendix B Analytic formulae for covariance matrices and mutual information